If log4 5=a and log5 6=b, then log3 2 is equal to


1
Expert's answer
2021-11-25T16:15:26-0500

"a\cdot b=\log_45\cdot \log_56=\log_46=\frac{1}{2}\log_26=\frac{1}{2}(\log_23+\log_22)=\frac{1}{2}(\log_23+1)"

"2ab=\log_23+1"

"\log_23=2ab-1"

"\log_3 2=\frac{1}{\log_2 3}=\frac{1}{2ab-1}"


Answer: "\log_32=\frac{1}{2ab-1}" .